OpenStudy (vampirediaries):

Identify the vertex, focus, and directrix of the parabola with the equation x^2 -6x-8y+49=0

OpenStudy (vampirediaries):

So far I have \[8y=(x-3)^{2} +49\] What do I do next?

OpenStudy (mathmate):

Ouch! Something is missing!

OpenStudy (vampirediaries):

Can you help me @mathmate

OpenStudy (mathmate):

The regrouping is missing a 9 somewhere. From \(x^2 -6x-8y+49=0\) we get \(x^2 -6x+49=8y\) \(8y=x^2 -6x+49\) \(8y=(x^2 -6x+9) -9 +49\) \(8y=(x-3)^2 +40\) \(8(y-5)=(x-3)^2\) This means that the vertex of the conic has been translated to (3,5). Can you take it from here?

OpenStudy (vampirediaries):

Why take away 9 from 49 to make it 40?

OpenStudy (mathmate):

Because we need to add a nine to complete the square of (x-3), so we need to subtract 9 on the same side of the equal sign in order not to change the value of the expression.

OpenStudy (vampirediaries):

Oooooh :) Thank you for explaining that

OpenStudy (mathmate):

Can you complete the rest of the question? If you want to confirm your answer, just post.

OpenStudy (vampirediaries):

How did you get the 5? \[8(y-5)=(x-3)^{2}\]

OpenStudy (mathmate):

Moved the 40 over to the left, and factor with 8. 8x-40=8(x-5)

OpenStudy (mathmate):

Sorry, I meant 8y-40=8(y-5)

OpenStudy (vampirediaries):

From the equation above, when written in standard form of a horizontal parabola \[x= \frac{ 1 }{ 4c }(y-k)^{2}+h\] ; Am I going to use the 8 in the equation?

OpenStudy (mathmate):

What you have is a parabola with a vertical axis, because y is linear, and x is to the second degree. You cannot arbitrarily change a parabola with a vertical axis to one with a horizontal axis. The equation \(8(y-5)=(x-3)^2\) means that h=3, k=5, and p=2 in the equation \((y-k)=\frac{(x-h)^2}{4p}\) where p is the distance of the focus from the vertex, and it is also the distance of the vertex from the directrix. (h,k) is the position of the vertex.

OpenStudy (vampirediaries):

P=2 ?

OpenStudy (vampirediaries):

@amistre64 @mathstudent55 @KayleeNeedsHelp can you help me?

OpenStudy (amistre64):

Identify the vertex, focus, and directrix of the parabola with the equation x^2 -6x-8y+49=0 x^2 -6x+9 -8y+49-9=0 (x-3)^2-8y+40=0 (x-3)^2 = 8y-40 (x-3)^2 = 8 (y-5) 1/8 (x-3)^2 = (y-5)

OpenStudy (amistre64):

vertex should be easily determined, the p parameter seems to be 2 yes: 4(2)=8

OpenStudy (amistre64):

the p parameter just tells us how far to go up and down from the vertex to find the focus and directrix

OpenStudy (vampirediaries):

How did you get 2 though or why 2? I don't remember from the equations I know of having the number 2 being multiplied by 4. Though \[a=\frac{ 1 }{ 4c }\] I just want to know why 4 is being multiplied by 2 to get 8.

OpenStudy (mathmate):

the standard equation is y=x^2/(4p) (some books say y=x^2/(4c) The equation was made in such a way to recognize the distance between the focus and the vertex, and also the distance between the vertex and the directrix. Note that if the sign of p (or c) is reversed, the parabola is also reversed. When p>0, the parabola opens upwards. 

OpenStudy (mathmate):

... A horizontal parabola opens to the right (positive x) if p>0.

OpenStudy (amistre64):

where does the 4 come from? its just a side effect of creating the general equation of a parabola to start with. we define a parabola as all the points (x,y) that are the same distance from a focus point (a,b). and a directrix line (x=x, y=k) \[\sqrt{(x-a)^2+(y-b)^2}=\sqrt{(x-x)^2+(y-k)^2}\] \[(x-a)^2+(y-b)^2=(x-x)^2+(y-k)^2\] \[(x-a)^2+(y-b)^2=(y-k)^2\] \[(x-a)^2+y^2-2by+b^2=y^2-2ky+k^2\] \[(x-a)^2+b^2-k^2=2by-2ky\] \[(x-a)^2+(b^2-k^2)=2y(b-k)\] now b-k is the distance between the focus and the directrix line, the parameter 'p' is half of (b-k). \[p=\frac{(b-k)}{2}\] \[2p=(b-k)\] \[4p=2(b-k)\] and this is why we get a generality with a 4p sitting in the mix. \[(x-a)^2+(b^2-k^2)=4p~y\]